# Thermodynamic Isomorphism of Transformers: A Lagrangian Approach to Attention Dynamics

### v1

@@ VI.1 Is Layer Depth Truly Time?
One might argue that Attention is a static operation. However, in deep models ($L\gg 1$), the representation $x(t)$ evolves continuously. In our Lagrangian, the time derivative $\dot{\rho}$ corresponds to the layer-wise rate of change of the attention distribution. If $\dot{\rho}\to 0$, then the system has reached a fixed point or equilibrium, meaning further layers do not alter the semantic interpretation. This aligns with the observation that deeper layers in Transformers often stabilize into a consistent attention pattern, which justifies our thermodynamic equilibrium analysis.
VI.2 Temperature vs. Variance Scaling
Some may suggest that $T=\sqrt{d_{k}}$ is merely a variance normalization trick. However, in statistical mechanics, temperature is defined as the inverse of the entropy-to-energy change response ($T^{-1}=\partial S/\partial U$). In high-dimensional spaces, the energy (dot product) variance naturally scales with $d_{k}$. Without renormalization by $\sqrt{d_{k}}$, the entropy $S$ would vanish (collapse to argmax) for any non-trivial energy difference. Thus, $\sqrt{d_{k}}$ plays the exact thermodynamic role of temperature: it maintains the system in a liquid state where information can flow and prevents it from reaching a frozen solid state (zero entropy).
VII Limitations and Future Directions
While our framework establishes a mathematically consistent mapping between Transformer architectures and physical systems, there are several fundamental limitations that require careful consideration and provide fertile ground for future investigation.
VII.1 The Nature of the Time Variable
Current Limitation. The most significant conceptual challenge in our formulation is interpreting the time variable $t$. We have adopted the convention, standard in the Neural ODE literature [22], of identifying $t$ with the continuous layer depth. Under this interpretation, $\dot{\rho}$ represents the rate of change of attention distribution across layers, and the Lagrangian dynamics describe the evolution of information states through the network architecture. However, this mapping introduces a subtle but important inconsistency. Within a single layer, the attention mechanism operates as an instantaneous, feed-forward computation rather than a genuine temporal evolution. The Softmax operation is computed in a single forward pass with no iterative dynamics. Our Lagrangian derivation implicitly assumes a quasi-static process, a sequence of infinitesimally slow equilibrium states. This is physically distinct from true dynamical evolution.
Distinction from Recurrent Dynamics. It is crucial to distinguish our framework from genuinely recurrent architectures where attention weights evolve iteratively. In standard Transformers, the “time evolution” exists only in the mathematical abstraction of layer-wise propagation.
Future Directions.
  • Continuous-depth Transformers: Formulate attention mechanisms explicitly in the Neural ODE framework with trainable continuous depth, allowing for genuine temporal dynamics within a single layer [22]. This would validate whether our Lagrangian correctly predicts the learned evolution operators.
  • Iterative Attention Refinement: Design architectures where the attention weights genuinely evolve through gradient descent on our proposed the free energy functional within each layer, testing whether this improves performance or interpretability.
  • Measurement Protocol: Develop rigorous operational definitions for measuring $\dot{\rho}$ empirically, for instance, by analyzing the layer-wise change in attention distributions under controlled perturbations.
VII.2 Thermodynamic Variables and Operational Definitions
Current Limitation. Our derivation of the equation of state relies on the information gas assumption, specifically, that the energy spectrum is self-averaging over the context window $V_{ctx}$. While this assumption is standard in statistical mechanics for large systems (thermodynamic limit), its validity in finite-sized attention layers with highly structured linguistic data requires empirical verification.
The Problem of Interaction. An ideal gas assumes non-interacting particles. However, tokens in a sentence have strong semantic correlations, or interactions. Therefore, a more accurate model might be a van der Waals gas for information, where an interaction term corrects the ideal pressure.
Future Directions.
  • Virial Expansion: Investigate deviations from the ideal law $\mathcal{P}V=T$ by measuring the compressibility factor $Z=\mathcal{P}V/T$ in real models. Deviations from unity would quantify the “semantic interaction strength” between tokens.
  • Independent Temperature Measurement: Instead of defining $T\equiv\sqrt{d_{k}}$, one could measure the “kinetic temperature” via the variance of the gradients or the attention logits, checking if the fluctuation-dissipation theorem holds.
VII.3 Experimental Validation and the Risk of Tautology
Current Limitation. The simulation presented in Appendix A demonstrates that the information efficiency constant $\kappa$ converges to a stable value as context length increases. However, this result is partially tautological. We defined $\kappa\equiv(\mathcal{P}\cdot V_{ctx})/(N\cdot T)$. Then we observed that it stabilized. This is analogous to defining a new constant by rearranging the ideal gas law and then confirming that it remains constant, a circular validation.
What Would Constitute Genuine Falsification? A rigorous test of our theory requires predictions that could fail in a measurable way. For example, if $\kappa$ is truly architecture-independent, then models with vastly different architectures should exhibit similar $\kappa$ values when normalized.
Future Directions.
  • Cross-Architecture Experiments: Measure $\kappa$ for Transformers, Mamba, RWKV, and hybrid architectures. If the framework is universal, $\kappa$ should be calculable from architectural primitives without fitting to data.
  • Thermodynamic Efficiency Benchmark: Define a “Carnot efficiency” for intelligence systems: $\eta_{AI}=\text{Useful Work (Accuracy)}/\text{Heat Input (FLOPs)}$. Test whether models optimized under our thermodynamic framework achieve higher $\eta_{AI}$ than standard architectures.
  • Interventional Experiments: Artificially manipulate $T$ (by scaling $\sqrt{d_{k}}$) during inference and measure whether the system’s entropy responds according to $S\propto\ln Z(T)$ as predicted by the partition function.
VII.4 The Goldstone Boson Hypothesis: Analogy or Mechanism?
Current Limitation. Our claim that rotary positional embeddings (RoPE) exploit massless Goldstone modes arising from spontaneous symmetry breaking should be understood as primarily analogical rather than mechanistic. While the underlying mathematical structure is suggestive, we have not yet rigorously demonstrated that the attention manifold undergoes a genuine phase transition characterized by a well-defined order parameter.
Future Directions.
  • Order Parameter Identification: Define a measurable order parameter (e.g., alignment of attention vectors in a learned subspace) and show that it exhibits spontaneous symmetry breaking during training.
  • Dispersion Relation: If Goldstone bosons truly exist, they should exhibit a characteristic $\omega\propto k$ dispersion relation (massless excitations). Testing the Fourier spectrum of attention patterns and testing this prediction could validate the hypothesis.
VII.5 Toward a Unified Physical Theory of Intelligence
The Broader Vision. This work is an initial step toward treating both artificial and biological intelligence, as physical phenomena governed by conservation laws, symmetries, and extremal principles. However, several grand challenges remain.
  • Learning Dynamics: Our current framework focuses on inference. A complete theory must describe the thermodynamics of gradient descent.
  • Non-equilibrium phenomena: Real AI systems operate far from equilibrium. Extending our framework to incorporate Jarzynski equality or maximum entropy production principles could yield new training algorithms.
  • Biological Neural Networks: The ultimate test of universality is whether our First Law of Information Dynamics applies to biological brains.
VIII Conclusion
In this paper, we have proposed a physics-inspired framework for artificial intelligence. By formalizing the Transformer architecture within the framework of Lagrangian dynamics and thermodynamics, we have demonstrated that:
@@ VIII Conclusion
    Softmax is the unique equilibrium state of a system that minimizes Helmholtz free energy.
@@
    The Query-Key interaction is isomorphic to the field-dipole interaction, creating an “alignment energy” landscape.
@@
    Learning and inference are distinct thermodynamic processes. The former is a chemical evolution ($\mu dN$) and the latter is a mechanical cycle ($PdV$).
  4. 4.
    Emergent phenomena such as scaling laws and grokking are phase transitions governed by the specific heat and the equation of state of the information gas.
We have proposed a framework, not a finished theory. The above-outlined limitations encourage the community to test, refine, or refute our core hypothesis: that the unreasonable effectiveness of Transformers reflects their adherence to fundamental physical principles. Whether this vision proves correct or serves merely as a useful fiction, the journey toward a physics-based understanding of intelligence has only just begun. Future work (Step 2) will extend this formalism to biological systems, analyzing the thermodynamic efficiency $\kappa$ of living connectomes.
Appendix A Theoretical Simulation of Information Efficiency
We conducted a theoretical simulation to validate the proposed equation of state for information gas (Eq. 22) and estimate the stability of the information efficiency constant ($\kappa$) under varying context window lengths ($V_{ctx}$).
Figure 1: Theoretical estimation of the value of $\kappa$. The plot shows the convergence of $\kappa$ for two different model architectures as a function of context window volume ($V_{ctx}$). The system exhibits a phase transition from an initial transient state (high variance) to a stable thermodynamic equilibrium, suggesting that $\kappa$ may be an intensive property of the intelligence system.
A.1 Simulation Setup and Results
We modeled two distinct Transformer configurations to test the universality of $\kappa$ (Model A: Small, Model B: Medium). The Information Density ($\mathcal{P}$) was derived from the inverse of the Perplexity ($PPL$) scaling law. The simulation revealed two distinct thermodynamic phases:
  1. 1.
    Transient phase ($V_{ctx}<100$): In the short context regime, the system exhibits high volatility. This corresponds to a non-equilibrium state in which the information gas has not yet filled the available volume.
  2. 2.
    Equilibrium phase ($V_{ctx}>200$): As the volume increases, $\kappa$ converges to a stable constant. This stability empirically validates the first law of information dynamics, demonstrating that for a fixed temperature ($T$) and particle number ($N$), the product of pressure and volume remains conserved relative to the system’s efficiency.

### v2

@@ VII.1 Is Layer Depth Truly Time?
One might argue that Attention is a static operation. However, in deep models ($L\gg 1$), the representation $x(t)$ evolves continuously. In our Lagrangian formulation, the time derivative $\dot{\rho}$ corresponds to the layer-wise rate of change of the attention distribution. If $\dot{\rho}\to 0$, the system has reached a fixed point or equilibrium. This aligns with the empirical observation that deeper layers in Transformers often stabilize into a consistent attention pattern, justifying our thermodynamic equilibrium analysis.
VII.2 The Dual Nature of Temperature
Critics often suggest that $T=\sqrt{d_{k}}$ is merely a variance normalization trick. However, we have shown that it plays the exact thermodynamic role of temperature ($T^{-1}=\partial S/\partial U$). Furthermore, our analysis of grokking reveals a crucial distinction:
  • Structural Temperature ($T_{struct}=\sqrt{d_{k}}$): Governs the entropy of the attention distribution during inference.
  • Dynamical Temperature ($T_{eff}\propto 1/\|W\|^{2}$): Governs the annealing process during training.
This duality explains why a model can be structurally frozen (converged weights) yet thermodynamically fluid (stochastic sampling) during generation.
VII.3 Critical Dynamics and Finite-Size Crossover
The most significant implication of our specific heat analysis is the proposal that grokking manifests as a thermodynamic crossover with critical-like characteristics. While our current experimental regime ($p\in[19,113]$) does not exhibit formal power-law divergence ($C_{v}\to\infty$), the consistent appearance of a robust fluctuation peak suggests that the system operates in a near-critical region. The system operates in a crossover regime between ordered memorization and disordered noise. The observed $C_{v}$ peak acts as a dynamical precursor to the representational reorganization, establishing it as a pseudo-critical phenomenon in finite-dimensional neural networks. This suggests that while strict universality classes might only emerge in the infinite-depth or infinite-width limits, the fundamental mechanics of intelligence are governed by the same fluctuation-dissipation principles found in traditional statistical physics.
VIII Limitations and Future Directions
While our framework establishes a mapping between Transformer architectures and physical systems, we acknowledge several critical limitations that define the scope of our current effective field theory.
VIII.1 The Nature of Time: Effective vs. Dynamical
Following the Neural ODE literature [22], we interpret layer depth as a continuous time variable $t$. However, our Lagrangian derivation implicitly assumes a quasi-static process, which is strictly valid only in the continuum limit ($L\to\infty$). In shallow architectures such as our 2-layer model, the discretization error may suppress the full development of the predicted attention dynamics. It is more physically grounded to interpret $t$ as an emergent effective time governing the renormalization group flow of information. Future work should investigate deeper architectures to determine if the discrete layers asymptotically approach the continuous action principle derived herein.
VIII.2 Mean-Field Approximation and Interaction
Our information gas model treats tokens as non-interacting particles, representing a mean-field approximation. In reality, natural language is a strongly correlated system with long-range dependencies. Consequently, our derived ideal gas-like equation of state holds primarily in the limit of low semantic density. We propose extending the framework by introducing a van der Waals-like equation of state: $(\mathcal{P}+a/V^{2})(V-b)=N\kappa T$, where $a$ represents semantic attraction (contextual correlation) and $b$ represents the exclusion volume of token embeddings (sparsity constraints).
VIII.3 Scaling and the Nature of the Crossover
Although we observed a robust $C_{v}$ peak consistently preceding the grokking transition, the peak magnitude did not exhibit asymptotic power-law divergence within the explored regime ($p\in[19,113]$). This suggests that in shallow Transformers, the transition behaves as a finite-size crossover rather than a formal thermodynamic phase transition.
Proposed Directions. To further investigate the potential for universal scaling:
@@ VIII.3 Scaling and the Nature of the Crossover
    Extended System Sizes: Perform experiments across at least two decades of system size ($p>1000$) to reach the potential asymptotic regime.
@@
    Depth-Scaling Analysis: Systematically increase layer depth ($L$) to test if the $C_{v}$ peak height begins to scale with system size as the continuum limit is approached.
@@
    Data Collapse: Investigate whether scaled fluctuation observables collapse onto a universal curve once the appropriate critical exponents are identified in deeper models.
VIII.4 Thermodynamic Interpretation of Hallucinations
We associated hallucinations with thermodynamic fluctuations ($TdS$). While thermal noise provides a minimal physical explanation for stochastic errors, hallucinations in large language models likely involve non-equilibrium mechanisms and dataset-driven biases. Future studies should incorporate non-equilibrium work relations (e.g., Jarzynski equality) to model the entropy production during the generation of out-of-distribution sequences.
IX Conclusion
We have developed an effective thermodynamic description of Transformer attention by formulating its dynamics on an information manifold and constructing a corresponding variational principle. Within a Shannon–Boltzmann entropy framework, the Softmax distribution appears as an equilibrium configuration minimizing a free energy functional, suggesting a formal bridge between attention mechanisms and canonical statistical mechanics.
At the macroscopic level, we introduced a fluctuation-based observable analogous to specific heat and showed empirically that its peak consistently precedes the grokking transition in modular arithmetic tasks. Although no asymptotic divergence is observed in shallow, finite systems, the reproducible enhancement of fluctuations indicates a critical-like crossover in the internal energy landscape. the results suggest that thermodynamic observables provide a quantitatively testable language for describing representational transitions in neural networks and motivate future investigations into scaling behavior in deeper architectures.